# PriPHiT: Privacy-Preserving Hierarchical Training of Deep Neural Networks

### v1

@@ Abstract
The training phase of deep neural networks requires substantial resources and as such is often performed on cloud servers. However, this raises privacy concerns when the training dataset contains sensitive content, e.g., face images. In this work, we propose a method to perform the training phase of a deep learning model on both an edge device and a cloud server that prevents sensitive content being transmitted to the cloud while retaining the desired information. The proposed privacy-preserving method uses adversarial early exits to suppress the sensitive content at the edge and transmits the task-relevant information to the cloud. This approach incorporates noise addition during the training phase to provide a differential privacy guarantee. We extensively test our method on different facial datasets with diverse face attributes using various deep learning architectures, showcasing its outstanding performance. We also demonstrate the effectiveness of privacy preservation through successful defenses against different white-box and deep reconstruction attacks.
@@ I Introduction
The following parts of the manuscript are structured as follows: In Section II, we discuss works that are related to the different parts of our proposed method. In Section III, we demonstrate our privacy-preserving training approach through the different phases of execution and discuss the attacks during the training phase. In Section IV, we show the results of extensive experiments on different facial attributes and the effectiveness of our method against different classification and reconstruction attacks in the training phase and testing phase. Finally, we conclude our work in Section V.
@@ III Proposed Method
Fig. 1 shows an overview of PriPHiT method and its modules, together with the steps of edge-cloud execution. It aims to train a deep architecture within an edge-cloud system to achieve accurate inference of the desired content $y$ from the input image $x$. This must be done in a way that the input image $x$ and its sensitive attribute $z$ are not shared with the cloud. Moreover, the sensitive attribute $z$ must not be extractable from the feature map that is shared with the cloud during the training or inference phases. Additionally, the input image $x$ must not be reconstructable from this shared feature map. This method is made of two steps: edge pre-training and edge-cloud training. In edge pre-training, an edge feature extractor is adversarially pre-trained using two early exits on the user’s training set images, the desired content labels and the sensitive content labels. The goal of this step is to make sure that the edge can preserve the user’s privacy and remove the sensitive content when it is connected to the cloud for edge-cloud training. In edge-cloud training, the edge is connected to the cloud and a high-performance cloud classifier (the cloud analyzer) is trained for the user’s task using the sent feature maps extracted from the training set images and the desired content labels. The adversarial training of the edge also continues using the early exit to make sure that privacy is preserved. Assuming a malicious cloud, in this step, the cloud also begins to attack the shared feature maps using sensitive content classification attackers and input reconstruction attackers. The details of these steps are described in the next parts.
@@ III-B Edge-Cloud Training
The cloud system is assumed to be untrusted. As a result, no raw input images $x$ or sensitive content labels $z$ are shared with it. However, we study a scenario that the untrusted cloud uses its own dataset to train attackers to extract the sensitive label or to reconstruct the original input. These attackers use a dataset that contains input images $x^{\prime}$ and sensitive labels $z^{\prime}$ to train themselves on the extracted feature maps coming from the edge. Its goal is to learn to classify sensitive content or reconstruct the original inputs from the received feature map. The training procedure is shown in Fig. 1-middle-left. In this study, we assume that the cloud classification attacker uses a similar architecture to the cloud analyzer and is trained with the same speed as the analyzer on the cloud. By doing this, after every epoch of classifier attacker training using its own dataset ($x^{\prime}$ and $z^{\prime}$), it can perform a simultaneous attack on the feature map that is extracted from the user private input $x$ (that is called $x_{E}$) and tries to perform a classification on its sensitive content. The simultaneous attack scenario during the edge-cloud training is shown in Fig. 1-middle-right. Algorithm 2 shows the training procedure of attackers and their simultaneous attack during the training on the user’s feature maps. In the next section, we assess the effectiveness of the proposed hierarchical training method through various experiments.
@@ IV-A Experimental Settings
Three popular architectures, VGG-11 [26], ResNet-18 [1], and MobileViT-xxs [27] are used to test the proposed framework. VGG-11 is divided between the edge and the cloud after the third convolution layer, ResNet-18 is divided after the third residual block, and MobileViT-xxs is divided after the third MobileNet-V2 Block ( Supplementary Material VI ). The edge analyzer and adversary are chosen to be tiny architectures made of one linear layer and one convolution layer. This convolution layer reduces the number of channels and compresses the signal which results in a reduction in the number of neurons needed in the fully connected layer. It helps to reduce the number of parameters in the fully connected layer and the whole edge side which reduces the memory footprint of the edge, making it deployable on low-memory edge devices. It is worth noticing that even though the edge analyzer and adversary are small, the experiments show their significant effectiveness. The number of parameters at the edge and the cloud for the proposed hierarchical training method is shown in Table I. According to the table, the number of parameters at the edge is orders of magnitude lower than the cloud. This makes it convenient to run the method on the available memory-constrained edge devices. The number of multiplication and accumulation operations (MACCs) for the training on the edge and on the cloud are also shown in Table II. Similarly, the number of MACCs in this method is lower in the edge in comparison to the cloud, that makes it convenient for slower edge devices. Additionally, according to Tables I and II, for both number of parameters and number of MACCs, the overhead of PriPHiT method (for the added modules such as early exits) is small. The only exception is the number of parameters in MobileViT-xxs architecture where half of the parameters belong to the PriPHiT modules, that is due to the tiny number of parameters that this architecture has in the edge as it uses compressed MobileNet-V2 blocks [28]. It is worth mentioning that the we showed the maximum percentage of PriPHiT overhead in Tables I and II that is only relevant in the few edge-pretraining epochs. These percentages reduce significantly in the edge-cloud training as the analyzer early exit in the edge is not needed anymore. We implemented our model using a powerful GeForce RTX 2080 Ti [29] with $11$ GB of memory (which represents the cloud) and a low-resource Quadro K620 [30] with $2$ GB of memory (which represents the edge). In the experiments, we consider that the untrusted cloud designs the reconstruction attacker based on the structure of the edge feature extractor layers according to the suggestions of [8] for inverted model designing. This is a realistic scenario since people often use well-known architectures and the untrusted cloud can guess the structure of the first layers at the edge, based on the last layers implemented in the cloud.
TABLE I: Comparison of the number of parameters as an indicator of memory footprint in the edge and the cloud for PriPHiT in different deep neural network architectures.
@@
TABLE II: Comparison of number of the training MACC operations in the edge and the cloud for PriPHiT in different deep neural network architectures.
@@
The edge pretraining is done for $n_{p}=5$ epochs. The number of pretraining epochs is selected to be low to comply with the resource-constrained edge device. Afterward, the edge-cloud training is done for $n_{t}=50$ epochs. At the same time, the classification and reconstruction attackers are trained for $50$ epochs and perform simultaneous attacks on the feature maps that are shared by the user with the cloud at each epoch. In the next step, the testing phase is done on an unseen subset of the dataset that evaluates the analyzer model, the classification attacker, and the reconstruction attacker. The privacy budget $\epsilon$ is set to $0.5$, $1$, and $2$ for the experiments and the threshold $T$ of the thresholding layer is set to $20$. For the VGG-11 architecture, the loss coefficient is $\lambda=6$ and the adversary update is done for $m_{a}=10$ iterations at every epoch. For the ResNet-18 and MobileViT-xxs architectures, we select $\lambda=9$ and $m_{a}=15$. To assess the effectiveness of the proposed privacy-preserving hierarchical training method, we propose a baseline that is also an edge-cloud training method aiming to achieve a high-accuracy classification of the desired content at the cloud using similar architectures (VGG-11, Resnet-18, and MobileNet-xxs are split at similar points). Note that the baseline does not use any privacy-preserving methods. Similar to our model, the edge-cloud training of the baseline is done for $50$ epochs and similar classification and reconstruction attackers are trained for $50$ epochs together with this model to execute simultaneous attacks throughout the training process. The baseline analyzer and attackers are also tested on the same unseen dataset at the end, during the inference phase.
@@ IV-B Training Phase and Simultaneous Attack
Figs. 2 and 3 show the accuracy of the simultaneous classification attack and the mean squared error between the original input samples and the reconstructions performed by the reconstruction attack during the training phase. The results are illustrated for VGG-11, ResNet-18, and MobileViT-xxs architectures with smiling as the desired content and gender as the sensitive content in the experiments on the CelebA dataset. In the baseline model, substantial privacy leakage occurs during the training process, as evidenced by the high-accuracy classification provided by the simultaneously trained classification attacker. However, in the PriPHiT method, the classification attackers fail to provide a high-accuracy classification of the sensitive content from the shared feature maps. Moreover, the trend reaches a plateau, indicating the inability to increase accuracy with more epochs due to the ongoing adversarial training leveraging the edge early exit adversary and the cloud analyzer. Regarding the simultaneous reconstruction attack during the training phase, the baselines exhibit low reconstruction loss with a decreasing trend across epochs, indicating a successful attack. However, in PriPHiT, the attack loss on the user’s shared feature map is high and does not change significantly with more epochs, which indicates an unsuccessful simultaneous attack due to the continuous privacy-preserving adversarial training of the feature extractor in the edge. Supplementary Material VII provides similar plots for experiments involving wearing makeup as the sensitive content and having an open mouth as the desired content. Additionally, Fig. 4 shows similar results for smiling versus gender experiments on the FFHQ dataset. Similar to the experiments on the CelebA dataset, we see that our proposed method reduces the accuracy of the simultaneous classification attacker significantly in comparison to the baselines. Additionally, although the average MSE losses of the reconstruction attacks are low for the baselines that indicate a successful attack, the reconstruction losses in the proposed PriPHiT method stay high, indicating unsuccessful attacks. VGG-11 and ResNet-18 architectures are used for the experiments on the FFHQ dataset. Due to the insufficient number of samples in this dataset, implementing and training transformer architectures is not feasible for this experiment.
@@
TABLE III: Comparison of the test accuracy of PriPHiT with the baselines for different privacy budgets using various architectures. The results are provided for the two experiments on the CelebA dataset: 1- Desired content: smiling, sensitive content: gender. 2- Desired content: having a slightly open mouth, sensitive content: wearing heavy makeup. For each group of experiments, a trivial classifier shows the lowest accuracy level. It is similar to a blind classifier that just outputs the class with the majority of labels.
Experiment Architecture Method Analyser Accuracy Attacker Accuracy
@@
TABLE IV: Comparison of the test accuracy of PriPHiT with the baselines for different privacy budgets using different architectures for the experiments on the FFHQ dataset. The results are provided when the task-relevant content is smiling and the sensitive content is selected to be gender. For the experiments, a trivial classifier shows the lowest accuracy level. It is similar to a blind classifier that just outputs the class with the majority of labels.
Experiment Architecture Method Analyser Accuracy Attacker Accuracy

### v2

@@ Abstract
The training phase of deep neural networks requires substantial resources and as such is often performed on cloud servers. However, this raises privacy concerns when the training dataset contains sensitive content, e.g., facial or medical images. In this work, we propose a method to perform the training phase of a deep learning model on both an edge device and a cloud server that prevents sensitive content being transmitted to the cloud while retaining the desired information. The proposed privacy-preserving method uses adversarial early exits to suppress the sensitive content at the edge and transmits the task-relevant information to the cloud. This approach incorporates noise addition during the training phase to provide a differential privacy guarantee. We extensively test our method on different facial and medical datasets with diverse attributes using various deep learning architectures, showcasing its outstanding performance. We also demonstrate the effectiveness of privacy preservation through successful defenses against different white-box, deep and GAN-based reconstruction attacks. This approach is designed for resource-constrained edge devices, ensuring minimal memory usage and computational overhead.
@@ I Introduction
The following parts of the manuscript are structured as follows: In Section II, we discuss works that are related to the different parts of our proposed method. In Section III, we demonstrate our privacy-preserving training approach through the different phases of execution and discuss the attacks during the training phase. In Section IV, we show the results of extensive experiments on different facial and medical datasets with various attributes and the effectiveness of our method against different classification and reconstruction attacks in the training phase and the inference phase. Finally, we conclude our work in Section V.
@@ II-B Preserving Privacy in the Training Phase
TABLE I: Summary of Key Contributions and Limitations in Related Works
Category Related Work                 Key Contributions                   Limitations
Private Feature Extraction Huang et al. [14], Li et al. [15] Adversarial training between edge feature extractor, desired classifier, and adversary to remove sensitive content while retaining desired information for the inference phase. No privacy guarantee for training phase; effectiveness against specific attacks only; lacks theoretical privacy guarantees.
Osia et al. [11] Uses an inverse Siamese network to push apart feature maps with similar sensitive labels and pull together those with different labels in the inference phase. Fails to address training phase privacy; no theoretical privacy guarantees provided.
Privacy in Training Phase Yu et al. [16] Homomorphic encryption for collaborative edge-cloud training; provides strong privacy guarantees. Resource-heavy for edge devices; applicable only to simple models and datasets.
Guo et al. [13] Noise added to feature maps for differential privacy; ensures strong privacy for transmitted data. Requires pre-trained feature extractors, limiting task-specific performance; sensitive information not explicitly suppressed.
Gupta and Raskar [20] Distributed training with split neural networks; encryption ensures parameter privacy. No protection for transmitted feature maps; requires sharing sensitive labels or multiple communication passes, increasing computational burden and communication costs.
Thapa et al. [21] Incorporates noise for differential privacy; collaborative training with two servers, one contains a part of the neural network architecture and one for the parameter averaging. Noise addition degrades task performance; requires sharing sensitive labels or multiple communication passes, increasing computational burden and communication costs.
PriPHiT Adversarial early exiting for selective sensitive content suppression; preserves the accuracy of the desired task; avoids label sharing and ensures robust defense against white-box and black-box attacks.                           —
Another category of related work focuses on Split Federated Learning (SFL), which is a technique for distributed training of neural networks by splitting models between clients and servers. Gupta and Raskar [20] introduced an SFL framework where a neural network is split between a cloud server and clients, enabling classification tasks. Their system encrypts and exchanges client parameters but lacks mechanisms to safeguard transmitted feature maps against reconstruction attacks. Thapa et al. [21] extended this work with Splitfed, involving multiple clients and two servers, and introduced noise-based differential privacy for client communicated parameters and feature maps. However, this approach leads to a substantial reduction in model accuracy, even on relatively simple tasks such as MNIST classification [24]. Both methods require sharing sensitive labels with the server, which is then mitigated by placing final network layers on the client side (in addition to the early layers), however, this substantially increases communication costs and computational demands.
In contrast, our work introduces a privacy-preserving hierarchical training method between an edge device and a cloud server, addressing limitations in existing approaches. By employing a novel adversarial early exiting technique during training, the method selectively retains desired information in transmitted feature maps to maintain task accuracy while eliminating sensitive content, even for complex tasks like facial attribute recognition where the desired information and the sensitive content are heavily mixed. Unlike prior methods, it robustly defends against white-box and advanced deep reconstruction attacks, without sharing sensitive labels or using workarounds that impose significant computational demands or communication costs on edge devices. Table I provides a summary of the related works discussed in this section, highlighting their limitations in comparison to our approach.
@@ III Proposed Method
TABLE II: List of Symbols Used in the Proposed Method
Symbol                        Description
$x$ Input image (e.g., face image).
$x^{\prime}$ Cloud’s input dataset for the attackers’ training.
$x_{E}$ Feature map extracted by the edge device from $x$.
$\overline{\overline{x^{\prime}}}$ Reconstruction attacker result in the attackers’ training.
$\overline{\overline{x}}$ Reconstruction attacker result in the simultaneous attack.
$\overline{\overline{x^{\prime\prime}}}$ Reconstruction attacker result in testing.
$y$ Desired content label in the training set.
$\tilde{y}$ Predicted desired content by the early-exit analyzer.
$\bar{y}$ Predicted desired content by the cloud analyzer in training.
$y^{\prime\prime}$ Desired content label in the test set.
$\overline{y^{\prime\prime}}$ Predicted desired content by the cloud analyzer in testing.
$z$ Sensitive content label in the training set.
$\tilde{z}$ Predicted sensitive content by the early-exit adversary.
$z^{\prime}$ Sensitive content label in the cloud’s attacker dataset.
$\overline{\overline{z^{\prime}}}$ Predicted sensitive content in the attackers’ training.
$\overline{\overline{z}}$ Predicted sensitive content in the simultaneous attack.
$\overline{\overline{z^{\prime\prime}}}$ Predicted sensitive content by the classification attacker in testing.
$\epsilon$ Privacy budget ensuring differential privacy.
$T$ Threshold to bound the feature map.
$m_{a}$ Number of adversarial training steps per epoch.
$n_{p}$ Number of pre-training epochs at the edge.
$n_{t}$ Number of edge-cloud training epochs.
@@
Fig. 1 shows an overview of PriPHiT method and its modules, together with the steps of edge-cloud execution. A complete list of the symbols used in the proposed method can also be found in Table II. Our method aims to train a deep architecture within an edge-cloud system to achieve accurate inference of the desired content $y$ from the input image $x$. This must be done in a way that the input image $x$ and its sensitive attribute $z$ are not shared with the cloud. Moreover, the sensitive attribute $z$ must not be extractable from the feature map that is shared with the cloud during the training or inference phases. Additionally, the input image $x$ must not be reconstructable from this shared feature map. This method is made of two steps: edge pre-training and edge-cloud training. In edge pre-training, an edge feature extractor is adversarially pre-trained using two early exits on the user’s training set images, the desired content labels and the sensitive content labels. The goal of this step is to make sure that the edge can preserve the user’s privacy and remove the sensitive content when it is connected to the cloud for edge-cloud training. In edge-cloud training, the edge is connected to the cloud and a high-performance cloud classifier (the cloud analyzer) is trained for the user’s task using the sent feature maps extracted from the training set images and the desired content labels. The adversarial training of the edge also continues using the early exit to make sure that privacy is preserved. Assuming a malicious cloud, in this step, the cloud also begins to attack the shared feature maps using sensitive content classification attackers and input reconstruction attackers. The details of these steps are described in the next parts.
@@ III-B Edge-Cloud Training
The cloud system is assumed to be untrusted. As a result, no raw input images $x$ or sensitive content labels $z$ are shared with it. However, we study a scenario that the untrusted cloud uses its own dataset to train attackers to extract the sensitive label or to reconstruct the original input. These attackers use a dataset that contains input images $x^{\prime}$ and sensitive labels $z^{\prime}$ to train themselves on the extracted feature maps coming from the edge. Its goal is to learn to classify sensitive content or reconstruct the original inputs from the received feature map. The training procedure is shown in Fig. 1-middle-left. In this study, we assume that the cloud classification attacker uses a similar architecture to the cloud analyzer and is trained with the same speed as the analyzer on the cloud. By doing this, after every epoch of classifier attacker training using its own dataset ($x^{\prime}$ and $z^{\prime}$), it can perform a simultaneous attack on the feature map that is extracted from the user private input $x$ (that is called $x_{E}$) and tries to perform a classification on its sensitive content. The simultaneous attack scenario during the edge-cloud training is shown in Fig. 1-middle-right. Algorithm 2 shows the training procedure of attackers and their simultaneous attack during the training on the user’s feature maps. A complexity analysis of the proposed training method is presented in Supplementary Material A, which demonstrates that the overall complexity is influenced by factors such as the computational speed of the edge device, the model dimensions, and the number of training samples and steps. Specifically, the edge pre-training phase dominates when the edge device has limited computational capacity or the number of edge pre-training epochs is high. Conversely, the edge-cloud training phase becomes the primary contributor to complexity if the cloud model is significantly large or the number of edge-cloud training epochs is substantial. Detailed runtime analysis and optimization, however, are beyond the scope of this study and are reserved for future work. In the next section, we assess the effectiveness of the proposed hierarchical training method through various experiments.
@@ IV-A Experimental Settings
TABLE III: Comparison of the number of parameters as an indicator of memory footprint in the edge and the cloud for PriPHiT in different deep neural network architectures.
@@
TABLE IV: Comparison of number of the training MACC operations in the edge and the cloud for PriPHiT in different deep neural network architectures.
@@
Three popular architectures, VGG-11 [30], ResNet-18 [1], and MobileViT-xxs [31] are used to test the proposed framework. VGG-11 is divided between the edge and the cloud after the third convolution layer, ResNet-18 is divided after the third residual block, and MobileViT-xxs is divided after the third MobileNet-V2 Block ( Supplementary Material B ). The edge analyzer and adversary are chosen to be tiny architectures made of one linear layer and one convolution layer. This convolution layer reduces the number of channels and compresses the signal which results in a reduction in the number of neurons needed in the fully connected layer. It helps to reduce the number of parameters in the fully connected layer and the whole edge side which reduces the memory footprint of the edge, making it deployable on low-memory edge devices. It is worth noticing that even though the edge analyzer and adversary are small, the experiments show their significant effectiveness. The number of parameters at the edge and the cloud for the proposed hierarchical training method is shown in Table III. According to the table, the number of parameters at the edge is orders of magnitude lower than the cloud. This makes it convenient to run the method on the available memory-constrained edge devices. The number of multiplication and accumulation operations (MACCs) for the training on the edge and on the cloud are also shown in Table IV. Similarly, the number of MACCs in this method is lower in the edge in comparison to the cloud, that makes it convenient for slower edge devices. Additionally, according to Tables III and IV, for both number of parameters and number of MACCs, the overhead of PriPHiT method (for the added modules such as early exits) is small. The only exception is the number of parameters in MobileViT-xxs architecture where half of the parameters belong to the PriPHiT modules, that is due to the tiny number of parameters that this architecture has in the edge as it uses compressed MobileNet-V2 blocks [32]. It is worth mentioning that the we showed the maximum percentage of PriPHiT overhead in Tables III and IV that is only relevant in the few edge pre-training epochs. These percentages reduce significantly in the edge-cloud training as the analyzer early exit in the edge is not needed anymore. We implemented our model using a powerful GeForce RTX 2080 Ti [33] with $11$ GB of memory (which represents the cloud) and a low-resource Quadro K620 [34] with $2$ GB of memory (which represents the edge). In the experiments, we consider that the untrusted cloud designs the reconstruction attacker based on the structure of the edge feature extractor layers according to the suggestions of [8] for inverted model designing. This is a realistic scenario since people often use well-known architectures and the untrusted cloud can guess the structure of the first layers at the edge, based on the last layers implemented in the cloud.
@@
The edge pre-training is done for $n_{p}=5$ epochs. The number of pre-training epochs is selected to be low to comply with the resource-constrained edge device. Afterward, the edge-cloud training is done for $n_{t}=50$ epochs. At the same time, the classification and reconstruction attackers are trained for $50$ epochs and perform simultaneous attacks on the feature maps that are shared by the user with the cloud at each epoch. In the next step, the testing phase is done on an unseen subset of the dataset that evaluates the analyzer model, the classification attacker, and the reconstruction attacker. The privacy budget $\epsilon$ is set to $0.5$, $1$, and $2$ for the experiments and the threshold $T$ of the thresholding layer is set to $20$. For the VGG-11 architecture, the loss coefficient is $\lambda=6$ and the adversary update is done for $m_{a}=10$ iterations at every epoch. For the ResNet-18 and MobileViT-xxs architectures, we select $\lambda=9$ and $m_{a}=15$. To assess the effectiveness of the proposed privacy-preserving hierarchical training method, we propose a baseline that is also an edge-cloud training method aiming to achieve a high-accuracy classification of the desired content at the cloud using similar architectures (VGG-11, Resnet-18, and MobileNet-xxs are split at similar points). Note that the baseline does not use any privacy-preserving methods. Similar to our model, the edge-cloud training of the baseline is done for $50$ epochs and similar classification and reconstruction attackers are trained for $50$ epochs together with this model to execute simultaneous attacks throughout the training process. The baseline analyzer and attackers are also tested on the same unseen dataset at the end, during the inference phase.
@@ IV-B Training Phase and Simultaneous Attack
Figs. 2 and 3 show the accuracy of the simultaneous classification attack and the mean squared error between the original input samples and the reconstructions performed by the reconstruction attack during the training phase. The results are illustrated for VGG-11, ResNet-18, and MobileViT-xxs architectures with smiling as the desired content and gender as the sensitive content in the experiments on the CelebA dataset. In the baseline model, substantial privacy leakage occurs during the training process, as evidenced by the high-accuracy classification provided by the simultaneously trained classification attacker. However, in the PriPHiT method, the classification attackers fail to provide a high-accuracy classification of the sensitive content from the shared feature maps. Moreover, the trend reaches a plateau, indicating the inability to increase accuracy with more epochs due to the ongoing adversarial training leveraging the edge early exit adversary and the cloud analyzer. Regarding the simultaneous reconstruction attack during the training phase, the baselines exhibit low reconstruction loss with a decreasing trend across epochs, indicating a successful attack. However, in PriPHiT, the attack loss on the user’s shared feature map is high and does not change significantly with more epochs, which indicates an unsuccessful simultaneous attack due to the continuous privacy-preserving adversarial training of the feature extractor in the edge. Supplementary Material C provides similar plots for experiments involving wearing makeup as the sensitive content and having an open mouth as the desired content. Additionally, Fig. 4 shows similar results for smiling versus gender experiments on the FFHQ dataset. Similar to the experiments on the CelebA dataset, we see that our proposed method reduces the accuracy of the simultaneous classification attacker significantly in comparison to the baselines. Additionally, although the average MSE losses of the reconstruction attacks are low for the baselines that indicate a successful attack, the reconstruction losses in the proposed PriPHiT method stay high, indicating unsuccessful attacks. VGG-11 and ResNet-18 architectures are used for the experiments on the FFHQ dataset. Due to the insufficient number of samples in this dataset, implementing and training the MobileViT-xxs architecture is not feasible for this experiment.
@@
TABLE V: Comparison of the test accuracy of PriPHiT with the baselines for different privacy budgets using various architectures. The results are provided for the two experiments on the CelebA dataset: 1- Desired content: smiling, sensitive content: gender. 2- Desired content: having a slightly open mouth, sensitive content: wearing heavy makeup. For each group of experiments, a trivial classifier shows the lowest accuracy level. It is similar to a blind classifier that just outputs the class with the majority of labels.
Experiment Architecture Method Analyzer Accuracy Attacker Accuracy
@@
TABLE VI: Comparison of the test accuracy of PriPHiT with the baselines for different privacy budgets using different architectures for the experiments on the FFHQ dataset. The results are provided when the task-relevant content is smiling and the sensitive content is selected to be gender. For the experiments, a trivial classifier shows the lowest accuracy level. It is similar to a blind classifier that just outputs the class with the majority of labels.
Experiment Architecture Method Analyzer Accuracy Attacker Accuracy